0.35g of an unknown compound dissolved in 100.0ml of water ahas an osmotic pressure of 32.6 mmhg at 15C. calculate the molar mass of the compound, answer 2 3 sig figs
Answer #1

π=mRT

= (W2*1000*RT)/(M2*w1)

M2=w2*1000*RT/π*w1

W2=0.35g, R=0.0821,T =288 ,

π =(32.6/760)=0.043 atm, w1=100g

=(0.35*1000*0.0821*288)/(0.043*100)

=1.92*10^3g

=1.92kg